OpenStudy (anonymous):

what is the limit as x approaches 0 when ((1-cosx)^2)/x

OpenStudy (amistre64):

0 i think

OpenStudy (amistre64):

2(1-cosx) . sin(x) = 0

OpenStudy (anonymous):

Yes L'pthals

OpenStudy (amistre64):

yep :)

OpenStudy (amistre64):

otherwise you expand the top and see what you can do to it..

OpenStudy (anonymous):

((1-cosx)(1-cosx))/x = 0?

OpenStudy (amistre64):

keep going....

OpenStudy (amistre64):

1 -2cos(X) + cos^2(x) -------------------- .... x

OpenStudy (anonymous):

doesn't (1-cosx)/x = o

OpenStudy (amistre64):

dunno; its been so long since I tried it the hard way :)

OpenStudy (amistre64):

you could try squeeze this on a graph

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=lim {x+to+0+}%28%281-cosx%29^2%29%2Fx

OpenStudy (anonymous):

But, (1-cosx)/x times (1-cosx)/x = (1-cosx)^2 / x^2.......

OpenStudy (anonymous):

Why would you do that?

OpenStudy (amistre64):

i dunno why it cuts it short; but copy paste that link itno your address bar

OpenStudy (anonymous):

@amogh - I'm just pointing out that I'm doing it wrong...

OpenStudy (anonymous):

Yes, x will go with either of them, not both of them

OpenStudy (anonymous):

@amistre64: do you like ff4?

OpenStudy (anonymous):

Sh@t! I did the problem wrong on my test then.......

OpenStudy (anonymous):

What did you answer?

OpenStudy (anonymous):

Thank you for the Wolfram alpha link amistre64. How much is that?

OpenStudy (anonymous):

I answered "0" but I didn't do my work correctly (analytically), so I get the problem wrong